a 4% salt solution is mixed with a 16% salt solution. How many millimeters of each solution are needed to obtain 600 milliliters of a 10% salt solution?
Answer by stanbon(75887)

a 4% salt solution is mixed with a 16% salt solution. How many millimeters of each solution are needed to obtain 600 milliliters of a 10% salt solution?
-----
Equation:
salt + salt = salt
0.04x + 0.16(600-x) = 0.10*600
-----
Multiply thru by 100 to get:
4x + 16*600 - 16x = 10*600
-12x = -6*600
x = 300 ml (amt. of 4% solution needed)
600-x = 300 ml (amt. of 16% solution needed)
===================
Cheers,
Stan H.
===================